Debt payments of $1,810.00 due today, $945 due in 113 days and $590.00 due in 325 days respectively are to be combined into a single payment to be made 160 days from now. Using an interest rate of 9.5% and the focal date of 160 days from now, determine the single payment.
Answer #1

Solution :-

Payment after 160 Days

Single Payment :-

Value of 1810 after 160 days = 1810*(1+9.5%*160/365) = 1810*(1 + 0.04164) = $1885.375

Value of 945 after 160 days that due in 113 days = 945*(1+9.5%*47/365) = 945*(1.01223) = $956.56

Value of 590 after 160 days that due in 325 days means 165 days = 590 / (1+0.095*165/325) = 590/1.0482 = $562.85

Single Payment = $1885.375 + $956.56 + $562.85 = $3404.789